### v1

# Convergence of integrals on the moduli spaces of curves and cographical matroids

Abstract.
We determine the convergence regions of certain local integrals on the moduli spaces of curves in neighborhoods of fixed stable curves in terms of the combinatorics of the corresponding graphs.
A.P. is partially supported by the NSF grants DMS-2001224, DMS-2349388, by the Simons Travel grant MPS-TSM-00002745, and within the framework of the HSE University Basic Research Program. N.P. is partially supported by the NSF grants DMS-1954050, DMS-2053243, and DMS-2344861.
1. Introduction
Let ¯$\overline{{\mathcal{M}}}_{g}$ denote the moduli space of stable curves of genus $g$, and let $\Delta^{ns}\subset\overline{{\mathcal{M}}}_{g}$ denote the non-separating node boundary divisor.
We consider a stratum ${\mathcal{M}}_{\Gamma}\subset\overline{{\mathcal{M}}}_{g}$, consisting of curves of given combinatorial type. Here $\Gamma$ is a stable (multi-)graph of genus $g$ (possibly with multiple edges and loops, with marking by genus on vertices). Let $\varphi$ be a top holomorphic form on ¯$\overline{{\mathcal{M}}}_{g}$ defined in a neighborhood in ¯$\overline{{\mathcal{M}}}_{g}$ of a point $C_{0}\in{\mathcal{M}}_{\Gamma}$, with a pole of order $1$ along $\Delta^{ns}$. We are interested in the region of convergence in $s\in\operatorname{\mathbb{C}}$ of the integral
$I_{\Gamma}(\varphi,s)=\int_{B_{\Gamma}}\frac{\varphi\wedge\overline{\varphi}}{|\det(\tau-\overline{\tau})|^{s}},$
where $B_{\Gamma}$ is a small ball in an étale neighborhood of $C_{0}$ in ¯$\overline{{\mathcal{M}}}_{g}$, and $\tau$ is the period matrix. Integrals of this type (for $s=5$) appear in calculation of vacuum amplitudes in superstring theory after integrating out the odd variables and using the GSO projection (see [3]).
The point is that $\det(\tau-\overline{\tau})$ has a logarithmic growth near $\Delta^{ns}$ that offsets the poles of $\varphi\wedge\overline{\varphi}$ for sufficiently large $s$. The precise region of convergence of $I_{\Gamma}(\varphi,s)$ depends on a graph $\Gamma$.
For example, if $\Gamma$ has a single vertex of genus $g-1$ with a loop, i.e., we are at the generic point of $\Delta^{ns}$, then $\det(\tau-\overline{\tau})=u\cdot\log|t|$, where $u$ is invertible and $t=0$ is a local equation of $\Delta^{ns}$. So the convergence of our integral for such $\Gamma$ is the same as for $\frac{dt\wedge d\overline{t}}{t\cdot\overline{t}\cdot(\log|t|)^{s}}$. Thus, in this case the integral absolutely converges for $\operatorname{Re}(s)>1$ and diverges for $s=1$.
Our main result, Theorem A below, determines the convergence threshold for each $\Gamma$. In particular, it shows that for each genus $\geq 6$, there exists a stable graph $\Gamma$ such that $I_{\Gamma}(\varphi,5)$ diverges. This means that the definition of superstring vacuum amplitudes for $g\geq 6$ requires some additional regularization procedure at the non-separating node boundary divisor in addition to the GSO projection.
Using the asymptotics of $\tau$ near $C_{0}$ (controled by the monodromy around branches of $\Delta^{ns}$), in the case when $C_{0}$ has rational components, we relate the integral $I_{\Gamma}$ to another integral defined in terms of the graph $\Gamma$.
Let $E(\Gamma)$ denote the set of edges of a connected graph $\Gamma$. We introduce independent variables $x_{e}$ associated $e\in E(\Gamma)$. Let $b=b(\Gamma)$ denote the 1st Betti number of $\Gamma$ (which is equal to $g$, the arithmetic genus of $C_{0}$), and let $c_{1},\ldots,c_{b}$ be a collection of simple cycles in $\Gamma$ giving a basis in $H_{1}(\Gamma)$. For each edge $e\in E(\Gamma)$, we define the index
$(c_{i},c_{j})_{e}=\begin{cases}1,&c_{i}\text{ and }c_{j}\text{ pass through }e\text{ in the same direction},\\ -1,&c_{i}\text{ and }c_{j}\text{ pass through }e\text{ in the opposite directions},\\ 0,&\text{ otherwise},\end{cases}$
and consider the $b\times b$-matrix $A=(a_{ij})$ with
$a_{ij}=\sum_{e\in E(\Gamma)}(c_{i},c_{j})_{e}\cdot x_{e}.$ (1.1)
Let $\psi_{\Gamma}\in\operatorname{\mathbb{Z}}[x_{e}]_{e\in E(\Gamma)}$ denote the Kirchoff polynomial (aka the first Symanzik polynomial) of $\Gamma$ (see [1, Sec. 2]), defined as the determinant
$\psi_{\Gamma}=\det(a_{ij}).$
This polynomial also has an expansion
$\psi_{\Gamma}=\sum_{T}\prod_{e\not\in T}x_{e},$ (1.2)
where $T$ runs over all spanning trees of $\Gamma$ (see [1, Prop. 3.4] and Prop. 2.1 below). Note that $\psi_{\Gamma}$ is homogeneous of degree $b(\Gamma)$.
Let $E^{\prime}(\Gamma)\subset E(\Gamma)$ denote the set of edges which are not bridges. We consider the integral
$J_{\Gamma}(s):=\int_{B}\frac{\prod_{e\in E(\Gamma)}dz_{e}d\overline{z}_{e}}{|\psi_{\Gamma}(\ln|z_{\bullet}|)|^{s}\cdot\prod_{e\in E^{\prime}(\Gamma)}z_{e}\overline{z}_{e}},$
where $B$ is a small ball around 00 in $\operatorname{\mathbb{C}}^{E(\Gamma)}$, $(z_{e})$ are complex coordinates on $\operatorname{\mathbb{C}}^{E(\Gamma)}$.
Given a connected graph $\Gamma$ without bridges (not necessarily stable), we define a rational constant $c(\Gamma)>0$ as follows. Consider the vector space $\operatorname{\mathbb{R}}^{E(\Gamma)}$ with the basis corresponding to edges of $\Gamma$, and let $v\in\operatorname{\mathbb{R}}^{E(\Gamma)}$ denote the sum of all basis vectors. For each spanning tree $T$, consider the vector
$v_{T}:=\sum_{e\not\in T}e\in\operatorname{\mathbb{R}}^{E(\Gamma)}.$
Now we set
$c(\Gamma)=\inf\left\{\sum c_{T}\ \big{|}\ \sum_{T}c_{T}v_{T}\geq v,c_{T}\in\operatorname{\mathbb{R}}_{\geq 0}\right\},$
where $w\geq v$ means that $w-v\in\operatorname{\mathbb{R}}_{\geq 0}^{E(\Gamma)}$.
It is easy to see that
$c(\Gamma)\geq\frac{e(\Gamma)}{b(\Gamma)},$ (1.3)
where $e(\Gamma)=|E(\Gamma)|$ is the number of edges and $b(\Gamma)$ is the 1st Betti number of $\Gamma$. Indeed, let $\varphi:\operatorname{\mathbb{R}}^{E(\Gamma)}\to\operatorname{\mathbb{R}}$ denote the map given by the sum of all coordinates. Then $\varphi(v)=e(\Gamma)$, $\varphi(v_{T})=b(\Gamma)$, so the inequality $\sum_{T}c_{T}v_{T}\geq v$ implies $b(\Gamma)\cdot(\sum_{T}c_{T})\geq e(\Gamma)$.
For an arbitrary connected $\Gamma$ (possibly with bridges), we set $c(\Gamma):=c(\Gamma^{\prime})$, where $\Gamma^{\prime}$ is obtained from $\Gamma$ by contracting all the bridges.
Theorem A. The integral $J_{\Gamma}(s)$ converges for $\operatorname{Re}(s)>c(\Gamma)$ and diverges for $s=c(\Gamma)$. If $\Gamma$ is stable and all components of $C_{0}$ are rational, the same assertions hold for the integral $I_{\Gamma}(\varphi,s)$.
A simple example is the $n$-gon graph $\Gamma=P_{n}$. It is easy to see that one has $c(P_{n})=n$. Indeed, the vectors $v_{T}$ are just the basis vectors $e_{i}$ and the condition $\sum c_{i}e_{i}\geq v$ means that $c_{i}\geq 1$, so the minimal $\sum c_{i}$ is $n$. There are similar stable graphs with all vertices of genus 00. For example, if $\Gamma$ is a $2n$-gon with every other side doubled then $c(\Gamma)=n$. Note that the genus of this graph is $n+1$. This shows that the boundary $\operatorname{Re}(s)=c$ of the convergence halfplane for $I_{\Gamma}(\varphi,s)$ can have arbitrary large $c$ as genus grows.
The proof of Theorem A uses some combinatorics of the cographical matroid associated to $\Gamma$ to reduce to the case of graphs $\Gamma$ for which inequality (1.3) becomes an equality (recall that the bases of the cographical matroid are complements to spanning trees, see [2, Sec. 2.3]). The key combinatorial result needed for Theorem A is that starting from any graph $\Gamma$, one can contract some edges in $\Gamma$ so that the resulting graph $\overline{\Gamma}$ satisfies $c(\overline{\Gamma})=e(\overline{\Gamma})/b(\overline{\Gamma})$ (see Cor. 3.7).
Convention. By a graph we mean a connected undirected multigraph, possibly with multiple edges and loops.
2. Kirchoff polynomial and the asymptotics of the period matrix
2.1. Kirchoff polynomial
We are going to give a proof of the determinant identity relating two definitions of $\psi_{\Gamma}$, using Cauchy-Binet theorem, and also get a bit more information about the corresponding matrix. There is also a simple recursive proof of this determinant formula in [1].
Let us consider a more general setup where we are given a surjective morphism of free $\operatorname{\mathbb{Z}}$-modules of finite rank
$\alpha:\operatorname{\mathbb{Z}}^{E}\to H,$
with the property that for every subset $S\subset E$, the cokernel $\operatorname{coker}(\alpha_{S})$ is a free $\operatorname{\mathbb{Z}}$-module, where $\alpha_{S}:\operatorname{\mathbb{Z}}^{S}\to H$ is the restriction of $A$.  [^1] Such a morphism is known as a unimodular collection of vectors in $H$.
We will apply it in the case when $E=E(\Gamma)$, $H=H^{1}(\Gamma)$, and $\alpha$ is the natural projection. Note that in this case $\operatorname{coker}(\alpha_{S})$ is canonically identified with $H^{1}(\Gamma^{S})$, where $\Gamma^{S}$ is obtained from $\Gamma$ by deleting edges in $S$.
With $\alpha$ as above we associate a symmetric bilinear form on $H^{*}$ with entries which are linear forms in independent variables $(x_{e})_{e\in E}$:
$B(\xi_{1},\xi_{2})=\sum_{e\in E}\xi_{1}(\alpha(e))\cdot\xi_{2}(\alpha(e))\cdot x_{e}.$
Thus, we have a well defined discriminant $\det(B)\in R:=\operatorname{\mathbb{Z}}[x_{e}\ |\ e\in E]$.
Let us say that $S\subset E$ is a basis if $\alpha_{S}:\operatorname{\mathbb{Z}}^{S}\to H$ is an isomorphism. We denote by ${\mathcal{B}}$ the set of all bases.
Proposition 2.1.
(i) One has $\det(B)=\sum_{S\in{\mathcal{B}}}\prod_{e\in S}x_{e}$.
(ii) Let us view $B$ as a nondegenerate form over the field $QR$, the fraction field over $R$, and let $B^{-1}$ denote the corresponding bilinear form on the dual space $H^{\vee}$ with values in $QR$. Then setting $x_{e}=\ln(y_{e})$, where $y_{e}>0$, we have
$\lim_{y\to 0}B^{-1}(\ln(y_{e}))=0.$
Proof.
(i) Let $f_{1},\ldots,f_{h}$ be a $\operatorname{\mathbb{Z}}$-basis of $H$, so we can view $\alpha$ as a matrix $E\times h$-matrix. Then the matrix of $B$ is given by
$B=\alpha\cdot D(x)\cdot\alpha^{t},$
where $D(x)$ is the diagonal $E\times E$-matrix with the entries $(x_{e})$. We can calculate $\det(B)$ by applying the Cauchy-Binet’s theorem to the decomposition of $B$ in $\alpha$ and $D(x)\alpha^{t}$. In this theorem we need to sum over subsets $S\subset E$, where $|S|=h$, such that the corresponding minor of $\alpha$ is nonzero. This condition is equivalent to the condition that $\alpha_{S}:\operatorname{\mathbb{Z}}^{S}\to H$ has nonzero determinant. Since $\operatorname{coker}(\alpha_{S})$ is free by assumption, this is equivalent to $S$ being a basis, in which case the corresponding minor is $\pm$. Since the corresponding minor of $\alpha^{t}$ is the same sign times $\prod_{e\in S}x_{e}$, the assertion follows.
(ii) In the dual basis $(f_{i}^{*})$ the matrix of $B^{-1}$ is the inverse matrix of $B$. Hence, it is enough to prove that every $(h-1)\times(h-1)$-minor $M$ of $B$ satisfies
$\lim_{y\to 0}\frac{M(\ln(y))}{\det B(\ln(y))}=0.$
Due to the formula for the determinant, it is enough to prove that every monomial appearing in $M$ is a constant multiple of $\prod_{e\in S}x_{e}$, for some $S\subset S^{\prime}$ with $|S|=h-1$ and $S^{\prime}$ a basis. Indeed, without loss of generality we can assume that $M$ corresponds to the rows $11,\ldots,h-$ 1. Then by the Cauchy-Binet’s formula, the monomials appearing in $M$ would correspond to some subsets $S\subset E$ with $|S|=h-1$ such that the map $\operatorname{\mathbb{Z}}^{S}\to H/\operatorname{\mathbb{Z}}\cdot f_{h}$ induced by $\alpha_{S}$ is nondegenerate. But this implies that $\alpha_{S}$ is injective, and so the image of $\alpha_{S}$ is of rank $h-1$. Since $\alpha$ is surjective, there exists an element $s\in E\setminus S$ such that $\alpha(s)$ is not contained in the image of $\alpha_{S}$, hence for $S^{\prime}=S\cup\{s\}$, the map $\alpha_{S^{\prime}}$ is nongenerate, i.e., $S^{\prime}$ is a basis. ∎
We are interested in the case of the natural projection $\alpha:\operatorname{\mathbb{Z}}^{E(\Gamma)}\to H^{1}(\Gamma)$ for a graph $\Gamma$. In this case, ${\mathcal{B}}$ is the set of bases of the cographical matroid associated with $\Gamma$. Choosing a basis of simple cycles $(c_{i})$, we can view the rows of $\alpha$ as coefficients of the edges in $c_{i}$ (with respect to a fixed orientation of all edges). Then the matrix of the symmetric form $B$ will be exactly the matrix (1.1). Hence, $\det(B)$ gets identified with the polynomial $\psi_{\Gamma}$ and we derive the expansion (1.2) Note that in this case bases are exactly complements to spanning trees. In particular, if an edge $e$ is a bridge then it is not contained in any bases, so $\psi_{\Gamma}$ depends only on variables corresponding to non-separating edges (i.e., edges that are not bridges).
2.2. Asymptotics of the period matrix
Assume that $\Gamma$ is a stable graph of genus $g$, with all vertices of genus 00, and let $C_{0}$ be a stable curve with rational components with the dual graph $\Gamma$. It is well known that the arithmetic genus of $C_{0}$ is $g$, so we can view $C_{0}$ as a point of ¯$\overline{{\mathcal{M}}}_{g}$. Then the set of edges $E(\Gamma)$ is in bijection with the branches of the boundary divisor through $C_{0}$, and the subset $E^{ns}(\Gamma)\subset E(\Gamma)$ of non-separating edges is in bijection with the branches of $\Delta^{ns}$ through $C_{0}$.
Note that if the graph $\Gamma$ is trivalent then the corresponding stratum ${\mathcal{M}}_{\Gamma}$ is a point and $e(\Gamma)=3g-3$. Otherwise, the stratum ${\mathcal{M}}_{\Gamma}$ has positive dimension.
For every non-separating edge $e$, we have the corresponding vanishing cycle $\alpha_{e}\in H_{1}(C)$, where $C$ is a smooth curve close to $C_{0}$, and the corresponding monodromy transformation $M_{e}$ on $H_{1}(C)$ has form
$M_{e}(x)=(x\cdot\alpha_{e})\alpha_{e}+x.$
More precisely, $C$ is obtained from the collection of spheres numbered by the set $V(\Gamma)$ of vertices of $\Gamma$ by connecting them with tubes numbered by the set of edges $E(\Gamma)$. We fix an orientation of $\Gamma$ and let $\beta_{e}$ denote a path along the tube corresponding to $e$ going in the direction of $e$. Then we define $\alpha_{e}$ as the class of a circle around the tube corresponding to $e$, so that $(\beta_{e}\cdot\alpha_{e})=1$.
The subgroup $A\subset H_{1}(C)$ generated by $(\alpha_{e})$ is maximal isotropic, and we have natural identifications
$H_{1}(C)/A\simeq H_{1}(C_{0})\simeq H_{1}(\Gamma)$
(see [1, Sec. 6]).
Let $(\alpha_{1},\ldots,\alpha_{g},\beta_{1},\ldots,\beta_{g})$ be a standard symplectic basis of $H_{1}(C)$ (so $(\beta_{j}\cdot\alpha_{i})=\delta_{ij}$) such that $A=\operatorname{\mathbb{Z}}\alpha_{1}+\ldots+\operatorname{\mathbb{Z}}\alpha_{g}$. In particular, $M_{e}$ does not change $\alpha_{i}$’s. Let $\omega_{1},\ldots,\omega_{g}$ be the basis of $H^{0}(C,\omega_{C})$, normalized by $\int_{\alpha_{i}}\omega_{j}=\delta_{ij}$. Then $M_{e}$ preserves $(\omega_{j})$, and changes the periods $\tau_{ij}=\int_{\beta_{i}}\omega_{j}$ to
$M_{e}:\tau_{ij}\mapsto\int_{M_{e}(\beta_{i})}\omega_{j}=\tau_{ij}+(\beta_{i}\cdot\alpha_{e})\cdot(\alpha_{e})_{j}=\tau_{ij}+(\beta_{i}\cdot\alpha_{e})\cdot(\beta_{j}\cdot\alpha_{e}),$
where the integers $(\alpha_{e})_{j}$ are determined from $\alpha_{e}=\sum_{j}(\alpha_{e})_{j}\cdot\alpha_{j}$.
Let $z_{e}$ denote a local equation of the branch of the boundary divisor near $C_{0}$ corresponding to the edge $e$. Then the monodromy $M_{e}$ acts on $\ln(z_{e})/(2\pi i)$ as $\ln(z_{e})/(2\pi i)\mapsto\ln(z_{e})/(2\pi i)+1$. Hence,
$\tau^{\prime}_{ij}:=\tau_{ij}-\sum_{e\in E}\frac{\ln(z_{e})}{2\pi i}\cdot(\beta_{i}\cdot\alpha_{e})\cdot(\beta_{j}\cdot\alpha_{e})$
are invariant under all monodromy transformations $M_{e}$. In other words, we have
$\tau=\tau_{0}(\frac{\ln(z)}{2\pi i})+\tau^{\prime},$ (2.1)
where $\tau=(\tau_{ij})$, $\tau^{\prime}=(\tau^{\prime}_{ij})$, and $\tau_{0}(x)$ is the matrix with coefficients in $\operatorname{\mathbb{Z}}[x_{e}\ |\ e\in E]$ with the entries
$\tau_{0}(x)_{ij}=\sum_{e\in E}x_{e}\cdot(\beta_{i}\cdot\alpha_{e})\cdot(\beta_{j}\cdot\alpha_{e}).$
In fact, it follows from the nilpotent orbit theorem (see [1, Sec. 9]) that in Eq. (2.1) the term $\tau^{\prime}$ is regular near $C_{0}$. We will use this to compute the asymptotics for $\det(\tau-\overline{\tau})$ near $C_{0}$.
Lemma 2.2.
Near $C_{0}$ one has
$\det(\tau-\overline{\tau})=\psi_{\Gamma}(\frac{\ln(|z|)}{\pi i})\cdot(1+f),$
where $z=(z_{e})$ and $f\to 0$ as $C\to C_{0}$.
Proof.
Let us choose a symplectic basis of $H_{1}(C)$ as follows. First, let $(c_{i})_{i=1,\ldots,g}$ be a basis of simple cycles in $H_{1}(\Gamma)$, and let $(\alpha_{i})_{i=1,\ldots,g}$ be the dual basis of $A$ with respect to the pairing between $A$ and $H_{1}(C)/A\simeq H_{1}(\Gamma)$ induced by the intersection pairing. Let $(\beta_{i})_{i=1,\ldots,g}$ be a set of mutually orthogonal classes in $H_{1}(C)$ projecting to $(c_{i})$ under the projection $H_{1}(C)\to H_{1}(\Gamma)$ (such a set exists since the intersection pairing is perfect). Then $(\alpha_{i},\beta_{i})$ is a standard symplectic basis. Furthermore, due to our definition of $\alpha_{e}$, the intersection index $(\beta_{i}\cdot\alpha_{e})$ is $1$ (resp., $-1$) exactly when $c_{i}$ passes through $e$ in the direction of the orientation (resp., in the opposite direction).
It follows that $\tau_{0}(x)$ coincides with the matrix $A=(a_{ij})$ given by (1.1). Since $\tau_{0}(x)$ depends on $(x_{e})$ linearly with integer coefficients, (2.1) gives
$\tau-\overline{\tau}=A(\frac{\ln|z|}{\pi i})+\tau^{\prime}-\overline{\tau^{\prime}}=A(\frac{\ln|z|}{\pi i})\cdot\bigl{(}1+A(\frac{\ln|z|}{\pi i})^{-1}\cdot(\tau^{\prime}-\overline{\tau^{\prime}})\bigr{)},$
where $\tau^{\prime}-\overline{\tau^{\prime}}$ is regular near $C_{0}$. Now by Proposition 2.1(ii), we have
$\lim_{C\to C_{0}}A(\frac{\ln|z|}{2\pi i})^{-1}\cdot(\tau^{\prime}-\overline{\tau^{\prime}})=0,$
and the assertion follows. ∎
3. Convergence region
3.1. Elementary observations
Lemma 3.1.
(i) The integral
$\int_{B}\frac{\prod_{i=1}^{n}dz_{i}d\overline{z}_{i}}{|\ln|z_{1}|\ldots\ln|z_{n}||^{s}\cdot z_{1}\overline{z}_{1}\ldots z_{n}\overline{z}_{n}},$
where $B$ is a small ball around the origin in $\operatorname{\mathbb{C}}^{n}$, converges for $s>1$ and diverges for $s=1$.
(ii) The integral
$\int_{B}\frac{\prod_{i=1}^{n}dz_{i}d\overline{z}_{i}}{|\ln|z_{1}|+\ldots+\ln|z_{n}||^{s}\cdot z_{1}\overline{z}_{1}\ldots z_{n}\overline{z}_{n}}$
converges for $s>n$ and diverges for $s=n$.
Proof.
(i) It is enough to consider the case $n=1$. Using polar coordinates, the assertion follows from to the fact that
$\int_{(0,\epsilon)}\frac{dr}{|\ln r|^{s}\cdot r}$
converges for $s>1$ and diverges for $s=1$.
(ii) Convergence for $s>n$ follows from the inequality $|\ln|z_{1}|+\ldots+\ln|z_{n}||^{n}\geq|\ln|z_{1}|\ldots\ln|z_{n}||$ together with part (i).
To prove the divergence for $s=n$, we will use induction on $n$. The base case $n=1$ follows from part (i). Now let $n>1$. Using polar coordinates and changing to the variables $y_{i}=\ln|z_{i}|$, the assertion is equivalent to the divergence of the integral
$\int_{y_{1}>C_{1}}\ldots\int_{y_{n}>C_{n}}\frac{dy_{1}\ldots dy_{n}}{(y_{1}+\ldots+y_{n})^{n}},$
for any large $C_{1},\ldots,C_{n}$. Performing the integration in $y_{n}$ we get
$\frac{1}{n-1}\int_{y_{1}>C_{1}}\ldots\int_{y_{n-1}>C_{n-1}}\frac{dy_{1}\ldots dy_{n-1}}{(y_{1}+\ldots+y_{n-1}+C_{n})^{n-1}}.$
It remains to use the change of variables $y_{1}\mapsto y_{1}+C_{n}$ and use the induction assumption. ∎
Lemma 3.2.
(i) The integral $J_{\Gamma}(s)$ converges for $s>c(G)$.
(ii) The integral $J_{\Gamma}(s)$ diverges for $s=e(\Gamma)/b(\Gamma)$.
Proof.
(i) Set $n=e(\Gamma)$. By Lemma 3.1(i), to prove convergence for $s>c(G)$ it is enough to show that
$\psi_{\Gamma}^{s}(x_{\bullet})\geq x_{1}\ldots x_{n}$
for $x_{i}\geq 1$. Indeed, we can assume $s=\sum c_{T}$, where $\sum c_{T}v_{T}\geq v$. Set $x_{T}=\prod_{e\not\in T}x_{e}$. Then for $x_{i}\geq 1$, we get
$\psi_{\Gamma}^{s}\geq\prod_{T}\psi_{\Gamma}^{c_{T}}\geq\prod_{T}x_{T}^{c_{T}}\geq x_{1}\ldots x_{n}.$
(ii) Set $b=b(\Gamma)$. Since $\psi_{\Gamma}$ has degree $b$, we have
$\psi_{\Gamma}\leq(x_{1}+\ldots+x_{n})^{b}.$
Now the divergence for $s=n/b$ follows immediately from Lemma 3.1(ii). ∎
Recall that we have an inequality $c(\Gamma)\geq e(\Gamma)/b(\Gamma)$ (see (1.3)).
Definition 3.3.
We say that $\Gamma$ optimal if $c(\Gamma)=e(\Gamma)/b(\Gamma)$.
Lemma 3.2 proves our assertion about convergence/divergence of $J_{\Gamma}(s)$ in the case of optimal $\Gamma$. Below we will reduce the case of a general $\Gamma$ to that of optimal $\Gamma$.
3.2. Combinatorial statement
Let $M$ be a loopless matroid on the ground set $E$. For all $S\subset E$, consider the linear functional
$\varphi_{S}:\operatorname{\mathbb{R}}^{E}\to\operatorname{\mathbb{R}}$
taking an element to the sum of its coordinates in $S$. The base polytope $P(M)$ consists of the vectors $w\in\operatorname{\mathbb{R}}^{E}$ such that $0\leq\varphi_{S}(w)\leq\operatorname{rk}S$ for all $S\subset E$ and $\varphi_{E}(w)=\operatorname{rk}E$.
Let $m:=\max\{|S|/\operatorname{rk}S\mid S\neq\emptyset\}$, and let $\mathcal{T}_{0}$ be the (nonempty) collection of subsets of $E$ that attain this maximum.
Lemma 3.4.
If $S,T\in\mathcal{T}_{0}$, then $S\cup T\in\mathcal{T}_{0}$. In other words, $\mathcal{T}_{0}$ has a maximal element.
Proof.
For any set $U\subset E$ of nonzero rank, we have $|U|\leq m\operatorname{rk}U$. Applying this inequality to $U=S\cap T$, we find that
$|S\cup T|$ $=$ $|S|+|T|-|S\cap T|$
$\geq$ $m\operatorname{rk}S+m\operatorname{rk}T-m\operatorname{rk}S\cap T$
$=$ $m(\operatorname{rk}S+\operatorname{rk}T-\operatorname{rk}S\cap T)$
$\geq$ $m\operatorname{rk}S\cup T.$
Applying it next to $U=S\cup T$, we find that $|S\cup T|=m\operatorname{rk}S\cup T$, thus $S\cup T\in\mathcal{T}_{0}$. ∎
Let
$c:=\min\{t\mid\text{there exists $w\in tP(M)$ with $w_{e}\geq 1$ for all $e\in E$}\},$
and let $w$ be an element of $cP(M)$ with $w_{e}\geq 1$ for all $e\in E$ (a witness for $c$).
Proposition 3.5.
We have $c=m$.
Proof.
Let $T_{0}$ be the maximal element of $\mathcal{T}$. Since $w_{e}\geq 1$ for all $e\in E$ and $w\in cP(M)$, we have $|T_{0}|\leq\varphi_{T_{0}}(w)\leq c\operatorname{rk}T_{0}$, and therefore $c\geq|T_{0}|/\operatorname{rk}T_{0}=m$.
Now we must prove the opposite inequality. We will do it by constructing an element $w\in mP(M)$ with $w_{e}\geq 1$ for all $e\in E$. This construction will proceed in stages.
First let $w(0)=(1,\ldots,1)$. We then have $\varphi_{S}(w(0))=|S|\leq m\operatorname{rk}S$ for all $S\subset E$, with equality if and only if $S\in\mathcal{T}_{0}$. In particular, we do not necessarily have $\varphi_{E}(w(0))=m\operatorname{rk}E$. If $T_{0}=E$, then $\varphi_{E}(w(0))=m\operatorname{rk}E$, so $w(0)\in mP(M)$ and we are done. If not, choose $e_{0}\in E\setminus T_{0}$. Then we have $e_{0}\notin S$ for all $S\in\mathcal{T}_{0}$. That means that there exists $\epsilon>0$ such that $\varphi_{S}(w(0)+\epsilon x_{e})\leq m\operatorname{rk}S$ for all $S\subset E$. Choose the largest such $\epsilon$, and let $w(1)=w(0)+\epsilon x_{e_{0}}$.
Now let $\mathcal{T}_{1}$ be the collection of subsets $S\subset E$ with the property that $\varphi_{S}(w(1))=m\operatorname{rk}S$. By an argument identical to that of Lemma 3.4, there is a maximal element $T_{1}$ of $\mathcal{T}_{1}$. If $T_{1}=E$, then $w(1)\in mP(M)$ and we are done. If not, choose $e_{1}\in E\setminus T_{1}$, and repeat the procedure to produce a new vector $w(2)$.
At some point we will have $T_{k}=E$, and this process will terminate with $w(k)\in mP(M)$ and $w(k)_{e}\geq 1$ for all $e\in E$. ∎
We’ll write $c(M)$ to denote the constant $c$ associated with a particular loopless matroid $M$.
Proposition 3.6.
Given a loopless matroid $M$, there exists a subset $T\subset E$ with the property that
$c(M)=c(M|_{T})=|T|/\operatorname{rk}T.$
Proof.
Let $T=T_{0}$ from the proof of Proposition 3.5. By the proposition, along with the fact that the rank function on $M|_{T}$ is the same as that on $M$, it is clear that $c(M)=c(M|_{T})=|T|/\operatorname{rk}T$. ∎
Corollary 3.7.
For any connected graph $\Gamma$ without bridges, there exists a graph $\overline{\Gamma}$ obtained by contracting some edges in $\Gamma$, such that
$c(\Gamma)=c(\overline{\Gamma})=\frac{e(\overline{\Gamma})}{b(\overline{\Gamma})}$
(the second equality means that $\overline{\Gamma}$ is optimal).
Proof.
Let $M$ denote the cographical matroid associated with $\Gamma$. Then $c(M)=c(\Gamma)$. The property that $\Gamma$ is optimal is the property that $c(M)=|E|/\operatorname{rk}E$. Corollary 3.6 says that we can find an optimal deletion of $M$ with the same constant $c$. That means that we can find an optimal contraction of $\Gamma$ with the same constant $c$. ∎
Example 3.8.
Let $\Gamma$ be an $2n$-gon with every other side doubled, so $e(\Gamma)=3n$. It is easy to see that $c(\Gamma)=n$. We can collapse all the doubled sides to get the $n$-gon $\overline{\Gamma}$, which is optimal and has $c(\overline{\Gamma})=n$.
3.3. Proof of Theorem A
Lemma 2.2 shows that the case of the integral $I_{\Gamma}(\varphi,s)$ (for stable $\Gamma$ and all components of $C_{0}$ rational) reduces to the case of $J_{\Gamma}(s)$. Also, since $\psi_{\Gamma}=\psi_{\Gamma^{\prime}}$, where $\Gamma^{\prime}$ is obtained by contracting all the bridges, it is enough to consider the case when $\Gamma$ has no bridges.
Due to Lemma 3.2(i), it remains to prove that $J_{\Gamma}(s)$ diverges for $s=c(\Gamma)$. Let $S\subset E(\Gamma)$ denote the set of edges that get contracted to get $\overline{\Gamma}$. By Fubini theorem, it is enough to prove that for any fixed values $z_{i}=c_{i}$ with $i\in S$, the integral
$\int_{B^{\prime}}\frac{\prod_{e\not\in S}dz_{e}d\overline{z}_{e}}{\psi_{\Gamma}(\ln|z_{\bullet}|)|_{z_{i}=c_{i},i\in S}^{c(\Gamma)}\cdot\prod_{e\not\in S}z_{e}\overline{z}_{e}},$
where $B^{\prime}$ is a small ball around the origin in $\operatorname{\mathbb{C}}^{E(\Gamma)\setminus S}$, diverges. Thus, by Lemma 3.1(ii) applied to $\overline{\Gamma}$, it is enough to prove an inequality
$\psi_{\Gamma}(x_{\bullet})|_{x_{e}=a_{e},e\in S}\leq C\cdot\psi_{\overline{\Gamma}}(x_{\bullet}),$
for $a_{e}>0$, with some constant $C>0$ depending on $(a_{e})$. Furthermore, it is enough that this inequality holds for $x_{e}>C$. Now we observe that for every spanning forest $T\subset\Gamma$, the intersection $\overline{T}:=T\cap\overline{\Gamma}$ is also a spanning forest. This implies that for every monomial $x_{T}=\prod_{\not\in T}x_{e}$ of $\psi_{\Gamma}$, one has
$x_{T}|_{x_{e}=a_{e},e\in S}=\prod_{e\in S}a_{e}\cdot x_{\overline{T}}.$
This easily leads to the claimed inequality.